OpenStudy (kaylak12345):

what is f(-5) if f(x)=l2x-1l+10

OpenStudy (saksh1998):

just put x = -5 mod would give out only positive value s0 21

OpenStudy (saksh1998):

fan and medal

OpenStudy (kaylak12345):

write an equation of the line with slope 2 that goes through point (3/2,6)

OpenStudy (saksh1998):

no need for that @kaylak12345 with all due respect

OpenStudy (francesca_marie):

|2x-1|=5 2x-1=-5 2x=-4 2x/2=-4/2 x=-2 answer is (x=3 and -2)

OpenStudy (kaylak12345):

um what question did you just answer there isnt any =-5

OpenStudy (francesca_marie):

I mis read it

OpenStudy (saksh1998):

we need to find the value of f(-5) so substitute x to be -5 in such questions thats the way

OpenStudy (kaylak12345):

what about the other question i asked

OpenStudy (saksh1998):